OpenStudy (anonymous):

If d/dx(f(3x^2))=9x^3, calculate f′(x)

jimthompson5910 (jim_thompson5910):

First I'm going to just focus on the left side Use the chain rule to get d/dx[f(3x^2)] = 9x^3 f ' (3x^2)*d/dx[3x^2] = 9x^3 f ' (3x^2)*6x = 9x^3 then divide both sides by 6x to get f ' (3x^2) = 1.5x^2 with me so far?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

Let g(x) = f ' (x) we know that g(3x^2) = 1.5x^2 what we want to find is g(x) ------------------------ Let's assume g(x) is a linear function. So g(x) = ax+b g(x) = ax+b g(3x^2) = a(3x^2)+b g(3x^2) = 3ax^2+b ---------------------- we know that g(3x^2) = 1.5x^2 and we found based on the linear assumption that g(3x^2) = 3ax^2+b, so let's equate the two 1.5x^2 = 3ax^2+b 1.5x^2+0 = 3ax^2+b we see that 1.5x^2 = 3ax^2 0 = b so we know b = 0 let's find the value of 'a' 1.5x^2 = 3ax^2 1.5 = 3a 3a = 1.5 a = 1.5/3 a = 0.5 So g(x) = 0.5x+0 = 0.5x This means f ' (x) = 0.5x

OpenStudy (anonymous):

thanks so much

jimthompson5910 (jim_thompson5910):

no problem